# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

### v1

@@ IV Conclusion
In summary, we have presented how to generate and manipulate squeezing-induced nonreciprocal quantum entanglement and asymmetric EPR steering in a three-mode COM system consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike. The WGM resonator supports two counterpropagating optical modes, which are uncoupled and degenerate. The coherent driving field from the waveguide is coupled into and out of the WGM resonator via evanescent coupling. Moreover, when placing the nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are orthogonal or parallel to the resonator surface. First, we show that when applying an additional two-photon parametric driving field for one specific input direction of the WGM resonator with a phase-matched squeezed vacuum reservoir, one can introduce a squeezed optical mode and an enhanced COM interaction in this input direction. This directional injection of quantum squeezing breaks the time-reversal symmetry of the system, and it allows one to tailor the parameter condition for entanglement generation with respect to its optical input direction, which is essential for achieving quantum nonreciprocity. Second, based on this unique feature of the system, we find that when reversing the input direction of the coherent driving field, quantum entanglement and the associated EPR steering is regulated in an asymmetric way, whereby a controllable nonreciprocal generation and manipulation of quantum entanglement and EPR steering is achieved. More interestingly, it is also found that by properly tuning the squeezing parameters, the directionality of EPR steering can be well controlled. According to this, we further show that with the increase of the squeezing strength, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. These results might open up new perspectives for the experimental generation and application of nonreciprocal bipartite and tripartite entanglement and asymmetric EPR steering as a precious resource for a variety of nascent quantum technologies ranging from one-sided device-independent quantum key distribution to no-cloning quantum teleportation Branciard et al. (2012); Walk et al. (2016); Li et al. (2024); He et al. (2015b); Chiu et al. (2016).
@@ V Acknowledgment
H.J. is supported by the National Natural Science Foundation of China (NSFC, Grant No. 11935006, 12421005), the National Key R&\&&D Program of China (Grants No. 2024YFE0102400), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the Science and Technology Innovation Program of Hunan Province (Grant No. 2020RC4047). L.-M.K. is supported by the NSFC (Grants No. 12247105, 11935006, 12175060 and 12421005), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the Henan Science and Technology Major Project (Grant No. 241100210400). Y.-F.J. is supported by the NSFC (Grant No. 12405029). D.-Y.W. is supported by the NSFC (Grant No. 12204424) and the China Postdoctoral Science Foundation (Grant No. 2022M722889). L.T. is supported by the NSFC (Grant No. 12305023) and the Sichuan Science and Technology Program (Grant No. 2024NSFSC1353). Y.W. is supported by the NSFC (Grant No. 12205256) and the Henan Provincial Science and Technology Research Project (Grant No. 232102221001). W.-S.B. is supported by the Henan Science and Technology Major Project of the Department of Science & Technology of Henan Province (Grant No. 241100210400).
@@ Appendix A Derivation of the effective Hamiltonian
To diagonalize the optical mode of Hamiltonian (1), one can perform a unitary Bogoliubov transformation with $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is a complex squeezing parameter with $r_{d}$ the squeezing strength and $\theta_{d}$ the squeezing reference angle. By implementing this Bogoliubov transformation, a squeezed optical mode is introduced and the system Hamiltonian (1) becomes
$\hat{\tilde{H}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\hat{H}_{\circlearrowleft}S(\eta_{d})$
$=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\chi\hat{q}_{1}\hat{q}_{2}$
@@
$\!-\!F_{1}\hat{q}_{1}\!-\!F_{2}\hat{q}_{2}\!+\!\chi\hat{q}_{1}\hat{q}_{2}\!+\!i\varepsilon_{d,\circlearrowleft}\!\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\!-\!\hat{a}_{s,\circlearrowleft})$
@@ Appendix B Derivation of the effective master equation under phase-matched condition
For the CCW input case, to eliminate the optical dissipation induced by the intracavity squeezing, we consider the injection of a broadband squeezed optical field in this direction, which can act as a squeezed-vacuum reservoir. We assume the squeezed optical field, with squeezing strength $r_{e}$ and reference phase angle $\theta_{e}$, is around the central frequency $\omega_{c}$. Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$. Then, including the dissipations of the optical and mechanical modes, the dynamics of the total system in the original picture is governed by the Born-Markovian master equation Gardiner and Zoller (2000)
$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{e}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
@@
and $\kappa~{}(\gamma_{m,j})$ is the optical (mechanical) decay rate. $\bar{n}_{m,j}=1/[\exp(\omega_{m,j}/k_{B}T)-1]$ is the thermal phonon number, while $N_{e}=\sinh^{2}(r_{e})$ and $M_{e}=\cosh(r_{e})\sinh(r_{e})e^{i\theta_{e}}$ describe the dissipation and the two-photon correlation of the cavity field caused by the squeezed-vacuum reservoir, respectively.
@@
$S^{\dagger}(\eta_{d})\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
@@
$S^{\dagger}(\eta_{d})\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}S(\eta_{d})$
@@
$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $S^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}S(\eta_{d})$
$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{s}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}+\frac{\kappa}{2}N_{s}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{s}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
@@
$N_{s}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e})+\cosh^{2}(r_{d})\sinh^{2}(r_{e})$
$+\dfrac{1}{2}\cos(\theta_{e}-\theta_{d})\sinh(2r_{d})\sinh(2r_{e}),$
$M_{s}=$ $e^{i\theta_{d}}[\cosh(r_{d})\cosh(r_{e})+e^{-i(\theta_{e}-\theta_{d})}\sinh(r_{d})\sinh(r_{e})]$
×$\times[\sinh(r_{d})\cosh(r_{e})+e^{i(\theta_{e}-\theta_{d})}\cosh(r_{d})\sinh(r_{e})].$ (30)
Here $N_{s}$ and $M_{s}$ denote the effective thermal noise and two-photon-correlation strength, respectively, which can be simplified in case of $r_{d}=r_{e}=r$, i.e.,
$N_{s}=$ $\dfrac{1}{2}\sinh^{2}(2r)[1+\cos(\theta_{e}-\theta_{d})],$
$M_{s}=$ $\dfrac{1}{2}e^{i\theta_{d}}\sinh(2r)[1+e^{i(\theta_{e}-\theta_{d})}]$
×$\times[\cosh^{2}(r)+e^{-i(\theta_{e}-\theta_{d})}\sinh^{2}(r)].$ (31)
Obviously, the thermal noise $N_{s}$ and the two-photon-correlation strength $M_{s}$ can be completely eliminated by choosing $r_{p}=r_{e}$ and $\theta_{e}-\theta_{d}=\pm n\pi(n=1,3,5...)$, i.e., $N_{s}=M_{s}=0$. As discussed in Ref. , the reservoir of the original optical mode is squeezed along the axis with angle $\dfrac{\theta_{e}}{2}$. In the basis of the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$, this effect can be canceled by the squeezing (along axis $\dfrac{\theta_{d}}{2}$) induced by the parametric amplification of $\hat{a}_{\circlearrowleft}$, when $r_{p}=r_{e}$ and $\theta_{e}+\theta_{d}=\pm n\pi(n=1,3,5...)$, yielding the phase-matched condition. This means the squeezed-vacuum reservoir of $\hat{a}_{\circlearrowleft}$ corresponds to an effective vacuum reservoir of $\hat{a}_{s,\circlearrowleft}$ under the phase-matched condition. In this case, the effective Born-Markovian master equation is derived as
@@ Appendix C Derivation of the dynamics of the quantum correlation function
$\dot{x}_{1}=-\kappa x_{1}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17},$ (34)
$\dot{x}_{2}=$ $-\kappa x_{2}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}-i\Lambda_{2}^{\ast}x_{17}+\kappa,$ (35)
$\dot{x}_{3}=2\omega_{m,1}x_{9}+2\omega_{m,1}x_{10},$ (36)
$\dot{x}_{4}=$ $-2\gamma_{m,1}x_{4}-2\omega_{m,1}x_{9}-2\omega_{m,1}x_{10}+2\Lambda_{1}x_{16}$
$+2\Lambda_{1}^{\ast}x_{15}-2\chi x_{24}+2\gamma_{m,1}\bar{n}_{m,1},$ (37)
$\dot{x}_{5}=2\omega_{m,2}x_{11}+2\omega_{m,2}x_{12},$ (38)
$\dot{x}_{6}=$ $-2\gamma_{m,2}x_{6}-2\omega_{m,2}x_{11}-2\omega_{m,2}x_{12}+2\Lambda_{2}x_{20}$
$+2\Lambda_{2}^{\ast}x_{19}-2\chi x_{23}+2\gamma_{m,2}\bar{n}_{m,2},$ (39)
$\dot{x}_{7}=-(2i\Delta_{s}+\kappa)x_{7}+2i\Lambda_{1}x_{13}+2i\Lambda_{2}x_{17},$ (40)
$\dot{x}_{8}=(2i\Delta_{s}-\kappa)x_{8}-2i\Lambda_{1}^{\ast}x_{14}-2i\Lambda_{2}^{\ast}x_{18},$ (41)
@@
$-\chi x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (45)
$\dot{x}_{13}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{13}+\omega_{m,1}x_{15}+i\Lambda_{1}x_{3}+i\Lambda_{2}x_{21},$ (46)
$\dot{x}_{14}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{14}+\omega_{m,1}x_{16}-i\Lambda_{1}^{\ast}x_{3}-i\Lambda_{2}^{\ast}x_{21},$ (47)
$\dot{x}_{15}=$ $-\left(i\Delta_{s}+\gamma_{m,1}+\dfrac{\kappa}{2}\right)x_{15}-\omega_{m,1}x_{13}+i\Lambda_{1}x_{10}$
$+i\Lambda_{2}x_{24}-\chi x_{17},$ (48)
$\dot{x}_{16}=$ $\left(i\Delta_{s}-\gamma_{m,1}-\dfrac{\kappa}{2}\right)x_{16}-\omega_{m,1}x_{14}-i\Lambda_{1}^{\ast}x_{10}$
$-i\Lambda_{2}^{\ast}x_{24}-\chi x_{18},$ (49)
$\dot{x}_{17}=-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{17}+\omega_{m,1}x_{19}+i\Lambda_{1}x_{21}+i\Lambda_{2}x_{5},$ (50)
$\dot{x}_{18}=\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{18}+\omega_{m,1}x_{20}-i\Lambda_{1}^{\ast}x_{21}-i\Lambda_{2}^{\ast}x_{5},$ (51)
@@
$\dot{x}_{22}=$ $-(\gamma_{m,1}+\gamma_{m,2})x_{22}-\omega_{m,1}x_{23}-\omega_{m,2}x_{24}-\chi x_{10}$
$-\chi x_{11}+\Lambda_{1}x_{20}+\Lambda_{1}^{\ast}x_{19}+\Lambda_{2}x_{16}+\Lambda_{2}^{\ast}x_{15},$ (55)
$\dot{x}_{23}=$ $-\gamma_{m,2}x_{23}+\omega_{m,1}x_{22}-\omega_{m,2}x_{21}-\chi x_{3}$
$+\Lambda_{2}x_{14}+\Lambda_{2}^{\ast}x_{13},$ (56)
$\dot{x}_{24}=$ $-\gamma_{m,1}x_{24}-\omega_{m,1}x_{21}+\omega_{m,2}x_{22}-\chi x_{5}$
$+\Lambda_{1}x_{18}+\Lambda_{1}^{\ast}x_{17}.$ (57)
@@
$\Omega_{3,\pm}$ $=i\Delta_{s}\pm\gamma_{m,2}\pm\dfrac{\kappa}{2}.,$ (59)
@@
$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (60)

### v2

@@ IV Conclusion
In summary, we have presented how to achieve phase-selective generation and manipulation of quantum entanglement and asymmetric EPR steering in a three-mode squeezed COM system. Here, as a specific example, we have considered a near-field cavity COM setup to realize our proposal [66, 67], which consists of a microdisk WGM resonator, a string of nanomechanical oscillators, and a tapered optical fiber. The WGM resonator supports two counterpropagating optical modes, which are degenerate and stay uncoupled. The coherent driving field from the waveguide is coupled into and out of the WGM resonator via evanescent coupling. Moreover, when placing the nanomechanical array close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field. First, we show that when simultaneously applying two-photon pumping fields and broadband squeezed optical fields to CW and CCW modes, one can introduce a pair of squeezed optical modes, with each coupled to an independent squeezed vacuum reservoir. This configuration enables a phase-controlled light-reservoir interaction and allows one to tailor asymmetric noise flows for opposite input directions, which plays a key role in breaking the symmetry of the quantum statistical properties of the counterpropagating modes. Second, based on this unique feature of the squeezed COM system, it is found that the generation of various types of quantum entanglement and the associated EPR steering become phase-dependent and thus they can be produced in an asymmetric way. More interestingly, it is also found that by properly tuning the intracavity squeezing parameters, the directionality of EPR steering can be well controlled. According to this, we further show that with the increase of the squeezing strength, the directionality of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. Note that our proposal can also be realized with various types of multimode COM system, e.g., microwave electromechanical system [23, 24], cavity magnomechanical system [76] , and nanospike-WGM optomechanical system [70]. These findings, opening up new perspectives for preparing rich types of quantum entanglement and EPR steering with asymmetric features, might be applicable for a variety of nascent quantum technologies ranging from one-sided device-independent quantum key distribution to no-cloning quantum teleportation [9, 10, 11, 12, 13].
@@ V Acknowledgment
H.J. is supported by the National Natural Science Foundation of China (NSFC, Grant No. 11935006, 12421005), the National Key R&\&&D Program of China (Grants No. 2024YFE0102400), and the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). L.-M.K. is supported by the NSFC (Grants No. 12247105, 11935006, 12175060 and 12421005), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the Henan Science and Technology Major Project (Grant No. 241100210400). Y.-F.J. is supported by the NSFC (Grant No. 12405029) and the Natural Science Foundation of Henan Province (Grant No. 252300421221). D.-Y.W. is supported by the NSFC (Grant No. 12204424) and the China Postdoctoral Science Foundation (Grant No. 2022M722889). L.T. is supported by the NSFC (Grant No. 12305023) and the Sichuan Science and Technology Program (Grant No. 2024NSFSC1353). Y.W. is supported by the NSFC (Grant No. 12205256) and the Henan Provincial Science and Technology Research Project (Grant No. 232102221001). W.-S.B. is supported by the Henan Science and Technology Major Project of the Department of Science & Technology of Henan Province (Grant No. 241100210400). Y.-L.Z. is supported by the Natural Science Foundation of Hunan Province (Grant No. 2025JJ60018) and the Scientific Research Foundation of Education Bureau of Hunan Province (Grant No.24B0866).
@@ Appendix A Derivation of the effective Hamiltonian
To diagonalize the optical mode of Hamiltonian (1), one can perform a unitary Bogoliubov transformation with $U(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is a complex squeezing parameter with $r_{d}$ the squeezing strength and $\theta_{d}$ the squeezing reference angle. By implementing this Bogoliubov transformation, a squeezed optical mode is introduced and the system Hamiltonian (1) becomes
$\hat{\tilde{H}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\hat{H}_{\circlearrowleft}U(\eta_{d})$
$=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\lambda\hat{q}_{1}\hat{q}_{2}$
@@
$\!-\!F_{1}\hat{q}_{1}\!-\!F_{2}\hat{q}_{2}\!+\!\lambda\hat{q}_{1}\hat{q}_{2}\!+\!i\varepsilon_{d,\circlearrowleft}\!\cosh(r_{d})(\hat{a}_{s,\circlearrowleft}^{\dagger}\!-\!\hat{a}_{s,\circlearrowleft})$
@@ Appendix B Derivation of the effective master equation under phase-matched condition
To regulate the light-reservoir interaction, we have considered the injection of a broadband squeezed optical field, which can act as a squeezed-vacuum reservoir. We assume the squeezed optical field, with squeezing strength $r_{e}$ and reference phase angle $\theta_{e}$, is around the central frequency $\omega_{c}$. Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$. Then, including the dissipations of the optical and mechanical modes, the dynamics of the total system in the original picture is governed by the Born-Markovian master equation [75]
$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e,\circlearrowleft}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{e,\circlearrowleft}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
@@
and $\kappa~{}(\gamma_{m,j})$ is the optical (mechanical) decay rate. $\bar{n}_{m,j}=1/[\exp(\omega_{m,j}/k_{B}T)-1]$ is the thermal phonon number, while $N_{e,\circlearrowleft}=\sinh^{2}(r_{e})$ and $M_{e,\circlearrowleft}=\cosh(r_{e})\sinh(r_{e})e^{i\theta_{e}}$ describe the dissipation and the two-photon correlation of the cavity field caused by the squeezed-vacuum reservoir, respectively.
@@
$U^{\dagger}(\eta_{d})\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}U(\eta_{d})$
@@
$U^{\dagger}(\eta_{d})\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}U(\eta_{d})$
@@
$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}U(\eta_{d})$
$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}+\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
@@
$N_{s,\circlearrowleft}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e,\circlearrowleft})+\cosh^{2}(r_{d})\sinh^{2}(r_{e,\circlearrowleft})$
$+\dfrac{1}{2}\cos\delta\theta\sinh(2r_{d})\sinh(2r_{e,\circlearrowleft}),$
$M_{s,\circlearrowleft}\!=$ $e^{i\theta_{d}}\![\sinh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{i\delta\theta_{\circlearrowleft}}\!\cosh(r_{d})\!\sinh(r_{e,\circlearrowleft})\!]$
×$\times\![\cosh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{-i\delta\theta_{\circlearrowleft}}\!\sinh(r_{d})\!\sinh(r_{e,\circlearrowleft})].$ (31)
Here $N_{s,\circlearrowleft}$ and $M_{s,\circlearrowleft}$ denote the effective thermal noise and two-photon-correlation strength, respectively, which can be simplified in case of $r_{d}=r_{e,\circlearrowleft}=r$, i.e.,
$N_{s,\circlearrowleft}=$ $\dfrac{1}{2}\sinh^{2}(2r)[1+\cos(\theta_{e,\circlearrowleft}-\theta_{d})],$
$M_{s,\circlearrowleft}=$ $\dfrac{1}{2}e^{i\theta_{d}}\sinh(2r)[1+e^{i(\theta_{e,\circlearrowleft}-\theta_{d})}]$
×$\times[\cosh^{2}(r)+e^{-i(\theta_{e,\circlearrowleft}-\theta_{d})}\sinh^{2}(r)].$ (32)
Obviously, the thermal noise $N_{s,\circlearrowleft}$ and the two-photon-correlation strength $M_{s,\circlearrowleft}$ can be completely eliminated by choosing $r_{d}=r_{e,\circlearrowleft}$ and $\theta_{e,\circlearrowleft}-\theta_{d}=\pm n\pi(n=1,3,5...)$, i.e., $N_{s,\circlearrowleft}=M_{s,\circlearrowleft}=0$. The reservoir of the original optical mode is squeezed along the axis with angle $\dfrac{\theta_{e,\circlearrowleft}}{2}$. In the basis of the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$, this effect can be canceled by the squeezing (along axis $\dfrac{\theta_{d}}{2}$) induced by the parametric amplification of $\hat{a}_{\circlearrowleft}$, when $r_{d}=r_{e,\circlearrowleft}$ and $\theta_{e,\circlearrowleft}+\theta_{d}=\pm n\pi(n=1,3,5...)$, yielding the phase-matched condition. This means the squeezed-vacuum reservoir of $\hat{a}_{\circlearrowleft}$ corresponds to an effective vacuum reservoir of $\hat{a}_{s,\circlearrowleft}$ under the phase-matched condition. In this case, the effective Born-Markovian master equation is derived as
@@ Appendix C Derivation of the dynamics of the quantum correlation function
$\dot{x}_{1}=$ $-\kappa x_{1}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}$
$-i\Lambda_{2}^{\ast}x_{17}+\kappa N_{s},$ (35)
$\dot{x}_{2}=$ $-\kappa x_{2}+i\Lambda_{1}x_{14}-i\Lambda_{1}^{\ast}x_{13}+i\Lambda_{2}x_{18}$
$-i\Lambda_{2}^{\ast}x_{17}+\kappa(N_{s}+1),$ (36)
$\dot{x}_{3}=\omega_{m,1}x_{9}+\omega_{m,1}x_{10},$ (37)
$\dot{x}_{4}=$ $-2\gamma_{m,1}x_{4}-\omega_{m,1}x_{9}-\omega_{m,1}x_{10}+2\Lambda_{1}x_{16}$
$+2\Lambda_{1}^{\ast}x_{15}-2\lambda x_{24}+2\gamma_{m,1}\bar{n}_{m,1},$ (38)
$\dot{x}_{5}=\omega_{m,2}x_{11}+\omega_{m,2}x_{12},$ (39)
$\dot{x}_{6}=$ $-2\gamma_{m,2}x_{6}-\omega_{m,2}x_{11}-\omega_{m,2}x_{12}+2\Lambda_{2}x_{20}$
$+2\Lambda_{2}^{\ast}x_{19}-2\lambda x_{23}+2\gamma_{m,2}\bar{n}_{m,2},$ (40)
$\dot{x}_{7}=-(2i\Delta_{s}+\kappa)x_{7}+2i\Lambda_{1}x_{13}+2i\Lambda_{2}x_{17}+\kappa M_{s}^{\ast},$ (41)
$\dot{x}_{8}=(2i\Delta_{s}-\kappa)x_{8}-2i\Lambda_{1}^{\ast}x_{14}-2i\Lambda_{2}^{\ast}x_{18}+\kappa M_{s},$ (42)
@@
$-\lambda x_{21}+\Lambda_{2}x_{18}+\Lambda_{2}^{\ast}x_{17},$ (46)
$\dot{x}_{13}=$ $-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{13}+\omega_{m,1}x_{15}$
$+i\Lambda_{1}x_{3}+i\Lambda_{2}x_{21},$ (47)
$\dot{x}_{14}=$ $\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{14}+\omega_{m,1}x_{16}$
$-i\Lambda_{1}^{\ast}x_{3}-i\Lambda_{2}^{\ast}x_{21},$ (48)
$\dot{x}_{15}=$ $-\left(i\Delta_{s}+\gamma_{m,1}+\dfrac{\kappa}{2}\right)x_{15}-\omega_{m,1}x_{13}+\Lambda_{1}^{\ast}x_{7}$
$+\Lambda_{1}x_{2}+i\Lambda_{1}x_{9}+i\Lambda_{2}x_{24}-\lambda x_{17},$ (49)
$\dot{x}_{16}=$ $\left(i\Delta_{s}-\gamma_{m,1}-\dfrac{\kappa}{2}\right)x_{16}-\omega_{m,1}x_{14}+\Lambda_{1}x_{8}$
$+\Lambda_{1}^{\ast}x_{1}-i\Lambda_{1}^{\ast}x_{9}-i\Lambda_{2}^{\ast}x_{24}-\lambda x_{18},$ (50)
$\dot{x}_{17}=$ $-\left(i\Delta_{s}+\dfrac{\kappa}{2}\right)x_{17}+\omega_{m,1}x_{19}$
$+i\Lambda_{1}x_{21}+i\Lambda_{2}x_{5},$ (51)
$\dot{x}_{18}=$ $\left(i\Delta_{s}-\dfrac{\kappa}{2}\right)x_{18}+\omega_{m,1}x_{20}$
$-i\Lambda_{1}^{\ast}x_{21}-i\Lambda_{2}^{\ast}x_{5},$ (52)
@@
$\dot{x}_{22}=$ $-(\gamma_{m,1}+\gamma_{m,2})x_{22}-\omega_{m,1}x_{23}-\omega_{m,2}x_{24}-\lambda x_{10}$
$-\lambda x_{11}+\Lambda_{1}x_{20}+\Lambda_{1}^{\ast}x_{19}+\Lambda_{2}x_{16}+\Lambda_{2}^{\ast}x_{15},$ (56)
$\dot{x}_{23}=$ $-\gamma_{m,2}x_{23}+\omega_{m,1}x_{22}-\omega_{m,2}x_{21}$
$-\lambda x_{3}+\Lambda_{2}x_{14}+\Lambda_{2}^{\ast}x_{13},$ (57)
$\dot{x}_{24}=$ $-\gamma_{m,1}x_{24}-\omega_{m,1}x_{21}+\omega_{m,2}x_{22}$
$-\lambda x_{5}+\Lambda_{1}x_{18}+\Lambda_{1}^{\ast}x_{17}.$ (58)
@@
$\Omega_{3,\pm}$ $=i\Delta_{s}\pm\gamma_{m,2}\pm\dfrac{\kappa}{2},$
$\Omega_{4}$ $=\gamma_{m,1}+\gamma_{m,2}.$ (60)
@@
$A=\left(\begin{matrix}A_{11}&A_{12}&A_{13}&A_{14}\\ A_{21}&A_{22}&A_{23}&A_{24}\\ A_{31}&A_{32}&A_{33}&A_{34}\\ A_{41}&A_{42}&A_{43}&A_{44}\\ \end{matrix}\right),$ (61)
$b=$ $\left[\kappa N_{s},\kappa(N_{s}+1),0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},\kappa M_{s}^{\ast},\right.$
$\left.\kappa M_{s},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (62)